The logistic growth model formula is given by
〖ye〗^kt+Ay=〖Le〗^kt
Where y I the population at time t(t≥0) and A,K and L are positive constants.
Use implicit differentiations to verify that
dy/dx=k/L y(L-y)

(d^2 y)/〖dt〗^2 =k^2/L^2 y(L-y)(L-2y)
1
Expert's answer
2020-06-14T17:39:56-0400

The logistic growth model formula is given


"ye^{kT}+Ay=Le^{kT}"

where "y" is the population at time "t (t\geq0)" and "A, k, L" are positive constants

Use implicit differentiations to verify that

(a)


"{dy\over dt}={k\over L}y(L-y)"

"{dy\over dt}e^{kT}+kye^{kT}+A{dy\over dt}=kLe^{kT}"


"{dy\over dt}(e^{kT}+A)=ke^{kT}(L-y)"

"e^{kT}+A={Le^{kT}\over y}"

Then


"{dy\over dt}={ke^{kT}(L-y)\over Le^{kT}}y"


"{dy\over dt}={k\over L}y(L-y)"

(b)


"{dy^2 \over dt^2 }={k^2 \over L^2 }y(L-y)(L-2y)"

"{dy^2 \over dt^2 }={d \over dt }({k\over L}y(L-y))={k\over L}({dy\over dt}(L-y)-y{dy\over dt})="


"={k\over L}{dy\over dt}(L-2y)={k\over L}({k\over L})y(L-y)(L-2y)="

"={k^2 \over L^2 }y(L-y)(L-2y)"